OpenStudy (anonymous):

solve the following exponential equation 9^x-14*3^x=-49 (a)What is the exact solution. (b)what is the decimal approximation for the solution.

OpenStudy (anonymous):

\[9^x-14(3^x)+49=0\\ 3^{2x}-14(3^x)+49=0\\ {\rm let}\quad u=3^x \]

jimthompson5910 (jim_thompson5910):

The equation is \[\Large 9^x - 14*3^x = -49\] right?

OpenStudy (anonymous):

becomes a quadratic

jimthompson5910 (jim_thompson5910):

if so, then use electrokid's method to get u^2 - 14u + 49 = 0 (u - 7)^2 = 0 u - 7 = 0 u = 7 3^x = 7

OpenStudy (anonymous):

now rewrite the equation in-terms of "u" variable

jimthompson5910 (jim_thompson5910):

keep going to solve for x

OpenStudy (anonymous):

right, so what you end up with is: \[3^x=7\] how do we bring the "x" down from the exponent?

OpenStudy (anonymous):

use the log and make it xlog3=7?

OpenStudy (anonymous):

you mean, \[x\log(3)=\log(7)\]

OpenStudy (anonymous):

yes!

jimthompson5910 (jim_thompson5910):

now divide both sides by log(3) to isolate x

OpenStudy (anonymous):

and voila!

OpenStudy (anonymous):

so the exact solution would be x= log of 7/ log of 3

OpenStudy (anonymous):

yep. you might also have .. \[x=\log_3(7)\]

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

**using change of base rule**

OpenStudy (anonymous):

I have a quick question would 4log(base5)(sqareroot of (9x-4)-log (base5) (2/x)+log(base 5) (2) would have the simplified form of log (base5)((9x-4)^2)/(2x-3))(2)

jimthompson5910 (jim_thompson5910):

where are you getting 2x-3 ?

OpenStudy (anonymous):

I really dont remember. but it feel like it was wrong. can you help me with it?

jimthompson5910 (jim_thompson5910):

ok each log is assumed to be base 5

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

4*log(sqrt(9x-4)) - log(2/x) + log(2) 4*log((9x-4)^(1/2)) - log(2/x) + log(2) log((9x-4)^((1/2)*4)) - log(2/x) + log(2) log((9x-4)^2) - log(2/x) + log(2) log((9x-4)^2) + log(2) - log(2/x) log(2(9x-4)^2) - log(2/x) log( [2(9x-4)^2] / (2/x) ) log( [2x(9x-4)^2] / 2 ) log( x(9x-4)^2 )

OpenStudy (anonymous):

thank super much! i can see the steps and understand it!

jimthompson5910 (jim_thompson5910):

ok glad it's making sense now